OpenStudy (anonymous):

Solve the equation : \[\frac{ b +2 }{3 }-\frac{ b -2 }{ 7}=2\] A. 5.1 B. 5.5 C. 6.0 D. 6.5

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\large \frac{ b +2 }{3 }-\frac{ b -2 }{ 7}=2\] \[\large 21*\frac{ b +2 }{3 }-21*\frac{ b -2 }{ 7}=21*2\] \[\large 7(b +2)-3(b -2)=42\] \[\large 7b +14-3b + 6=42\] I'll let you finish

OpenStudy (anonymous):

im pretty sure the answer is A

jimthompson5910 (jim_thompson5910):

no it's not, but it's close

OpenStudy (anonymous):

So it B ? it cause when i place them in its doesnt add to 42 :(

jimthompson5910 (jim_thompson5910):

yeah it's 5.5